OpenStudy (anonymous):

how do i do this: write an equation in slope intercept form as the line described; has the same slope s 3x-y+4=0 and contains the point(2,-3) ??

OpenStudy (anonymous):

i dont get it tho :(

OpenStudy (anonymous):

i dont know how!!

OpenStudy (superdavesuper):

general line eqn: ax+by+c=0 If two lines are parallel, their a n b will be the same so how would the new equation looks like?

OpenStudy (anonymous):

i have no idea cuz i dont even know whats the 'b' in this problem!

OpenStudy (superdavesuper):

Why not? The question says the new line is parallel to the line given right?

OpenStudy (superdavesuper):

"general line eqn: ax+by+c=0 If two lines are parallel, their a AND b will be the same"

OpenStudy (anonymous):

no it doesnt say anything about the new line being parallel!

OpenStudy (superdavesuper):

OK sorry my fault - parallel means has the same slope :)

OpenStudy (anonymous):

ohhhhh okay let me c now. hold on

OpenStudy (anonymous):

ummm i still have NO IDEA what to do

OpenStudy (superdavesuper):

Read what I said above and give me a while guess on how the new eqn will look like plz?

OpenStudy (anonymous):

wat am i supposed to do exactly tho? plug in numbers or wat?

OpenStudy (superdavesuper):

The plan is to use the given condition of the new line has the same slope to write out the new formula; then plug in the numbers given to find c

OpenStudy (anonymous):

i did plug the numbers in and i came up w/13=0. tht makes NO SENSE.

OpenStudy (superdavesuper):

no no no...ok lets say new eqn is ax+by+c=0 then it has same slope as 3x-y+4=0 with me so far?

OpenStudy (anonymous):

yes

OpenStudy (superdavesuper):

Two lines with same slope have the same a and b so....

OpenStudy (anonymous):

wats a and b?

OpenStudy (superdavesuper):

a and b are constants in the eqn ax+by+c=0

OpenStudy (anonymous):

is ax+by+c=0 the equation for finding an equation that has the same slope as this problem?

OpenStudy (superdavesuper):

ax+by+c=0 is the new equation the question is asking for but u need to find a, b and c.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

now what else do i do?

OpenStudy (superdavesuper):

Two lines with same slope have the same a and b

OpenStudy (anonymous):

okay

OpenStudy (superdavesuper):

new eqn is ax+by+c=0 then it has same slope as 3x-y+4=0 so what are a and b in this case?

OpenStudy (anonymous):

ummm i dont know

OpenStudy (superdavesuper):

Please read what I have just said in my two posts above.

OpenStudy (anonymous):

um im guessing tht a and b r the coordinates for a point

OpenStudy (superdavesuper):

x+y+4=0 is an eqn; so is 3x-5y-9=0 or -x+4y-2=0 they all have different a and b....first one a=1, b=1. second one a=3, b=-5. last a=-1, b=4.

OpenStudy (anonymous):

oh i get it! so wat do i do after?

OpenStudy (superdavesuper):

new eqn is ax+by+c=0 then it has same slope as 3x-y+4=0 so what are a and b in this case?

OpenStudy (anonymous):

theyre the constant right?

OpenStudy (superdavesuper):

YES :)

OpenStudy (anonymous):

and then...

OpenStudy (superdavesuper):

Show me what u think the new eqn looks like plz?

OpenStudy (anonymous):

3(2) -(-3) +4=0???

OpenStudy (superdavesuper):

How do u know the constant c? i only said a and b are the same...

OpenStudy (anonymous):

oh never mind it would be 3(2)-(-3)+4=0 right?

OpenStudy (superdavesuper):

Please read "How do u know the constant c? i only said a and b are the same..."

OpenStudy (anonymous):

cuz its given to me. they already put 4 as the number

OpenStudy (superdavesuper):

But thats the old eqn, new eqn the Q is asking is ax+bx+c=0 and u find a and b because it has the same slope as 3x-y+4=0. We do not know what c is....

OpenStudy (superdavesuper):

Do u understand that?

OpenStudy (anonymous):

would it be the same thing i told u b4 without the +4 tho?

OpenStudy (superdavesuper):

It is not the same because it is a different line eqn >.<

OpenStudy (anonymous):

i have no idea wat u r talking about, srry

OpenStudy (superdavesuper):

Will u take me up on my word that u dont know the value of c and write out in x, y term what the new eqn is plz?

OpenStudy (anonymous):

x-y=0

OpenStudy (anonymous):

um i think it would be better if i asked my teacher tomorow in school. thanks though!!

OpenStudy (superdavesuper):

Sorry I tried :(